Obtain the equation of the conic, a focus of which lies at (2,1), the directrix of which is x+y = 0 and which passes through (1,4). Also identify the conic.


1
Expert's answer
2021-06-01T18:18:47-0400

Answer:-

The distance of point (1,4)  from focus at (2,1) is

"\sqrt{
(
1
−
2
)^2

+
(
4
−
1
)^2
}
=
\sqrt{
10}"

The distance of point(1,4) from directrix is "\frac{1+4}{\sqrt{2}}=\frac{5}{\sqrt{2}}"

and ratio of distances is "\frac{\sqrt{10}}{\frac{5}{\sqrt{2}}}=\frac{\sqrt{20}}{5}=\frac{2}{\sqrt{5}}<1"

As the ratio is less than  1, it is an ellipse

and the equation is obtained from ratio of the distance of a point on ellipse say (x,y) from focus (2,1) and its distance from directrix 

x+y=0 being "\frac{2}{\sqrt{5}}" . the latter is "\frac{x+y}{\sqrt{2}}" , hence equation is

"\frac{(x-2)^2+(y-1)^2}{(\frac{x+y}{\sqrt{2}})^2}" ="(\frac{2}{\sqrt{5}})^2=\frac{4}{5}"

"\implies \boxed{2x^2-4xy+3y^2-20-10y+25=0}"